OpenStudy (anonymous):

can someone help me with partial derivatives? y= (LK)/(K+2L) derivative of Y in respect to L and derivative of Y in respect to K

myininaya (myininaya):

So when you take the partial of y w.r.t L you treat everything like a constant except L (and well y)

OpenStudy (anonymous):

yeah i know that

myininaya (myininaya):

You also need to apply quotient rule since you have a quotient.

myininaya (myininaya):

\[y_L=\frac{(LK)_L \cdot (K+2L) - (LK) \cdot (K+2L)_L}{(K+2L)^2}\]

OpenStudy (anonymous):

let me go straight to the pointi have to get this \[\frac{ derivative y respect \to L }{ derivative of Y respect \to K }\] which is supposed to be \[\frac{ K^2 }{ 2L^2 }\] but i'm getting instead \[\frac{ 2L^2 }{ K^2 }\]

myininaya (myininaya):

Did you apply quotient rule? I actually already applied the quotient rule and used little symbols there or notated what needed to be taken derivative of. Right now I'm not seeing how we are to got the 2L^2 on bottom.

OpenStudy (phi):

the good news is that the book is correct: K^2 / (2 L^2)

OpenStudy (anonymous):

yeah, i know... but that's what i'm getting in the book. i'm doing math for economics Y is the production function and i have to find the MRTS

OpenStudy (anonymous):

so, why is the book correct?

OpenStudy (phi):

what do you get for dy/dL ?

OpenStudy (anonymous):

4KL +4L^2

OpenStudy (phi):

use Mya's post

OpenStudy (anonymous):

dy/dk =\[k^2+4KL + 4L^2\]

OpenStudy (anonymous):

i'm still getting the same thing T_T

OpenStudy (phi):

can you do just one step? using the quotient rule ?

OpenStudy (phi):

ie. what is the numerator of dy/dL

OpenStudy (phi):

don't bother simplifying.

OpenStudy (anonymous):

dy/dl= \[\frac{ K.(K+2L)-2.(LK) }{ K^2+4KL+4L^2 }= \frac{K^2+2KL-2KL }{ K^2+4KL+4L^2 }= 4KL +4L^2\]

OpenStudy (anonymous):

ahh too late

OpenStudy (phi):

the top simplifies to k^2

myininaya (myininaya):

\[\frac{5}{5+9} \neq 9\]

OpenStudy (phi):

and I would leave the bottom as (k+2L)^2 (because it will cancel with the bottom from dy/dK)

myininaya (myininaya):

It looks like you canceled the some addends from top and bottom and left the bottom as the top

myininaya (myininaya):

You can cancel common factors

ganeshie8 (ganeshie8):

\[\large \dfrac{k^2}{k^2+4KL+4L^2} \ne \dfrac{1}{4KL+4L^2} \ne 4KL+4L^2\]

myininaya (myininaya):

but not common addends

OpenStudy (phi):

yes, whatever you did to simplify is way wrong. Better debug that. in the mean time you should just get \[ \frac{K^2}{(K+2L)^2 } \]

OpenStudy (phi):

I assume you notice that in the top 2KL - 2KL is 0

OpenStudy (anonymous):

oh waw that was a bad mistake

OpenStudy (anonymous):

thank you guys, i was really confused, i always mix the cancellation laws

OpenStudy (phi):

The "laws" also work for numbers. so if you try a similar problem with numbers, and you do something wrong, it will be obvious it did not work.